OpenStudy (anonymous):

Solve the system of equations algebraically & state its solution. a/2 + b/3= 3 a/3 + b/2= 2

OpenStudy (anonymous):

a = 2(3 - b/3) = 6 - 2/3 b

OpenStudy (anonymous):

what's next?

OpenStudy (anonymous):

let's get rid of the fractions, multiply by 6 3a + 2b = 18 2a + 3b = 12

OpenStudy (anonymous):

multiply eqn 2 by 3/2 3a + 9/2 b = 18

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

Okay I just did it. a=6 and b=0

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Alright thank you (: